# Solving Inverse Problems in Protein Space Using Diffusion-Based Priors

### v1

@@ Abstract
The interaction of a protein with its environment can be understood and controlled via its 3D structure. Experimental methods for protein structure determination, such as X-ray crystallography or cryogenic electron microscopy, shed light on biological processes but introduce challenging inverse problems. Learning-based approaches have emerged as accurate and efficient methods to solve these inverse problems for 3D structure determination, but are specialized for a predefined type of measurement. Here, we introduce a versatile framework to turn raw biophysical measurements of varying types into 3D atomic models. Our method combines a physics-based forward model of the measurement process with a pretrained generative model providing a task-agnostic, data-driven prior. Our method outperforms posterior sampling baselines on both linear and non-linear inverse problems. In particular, it is the first diffusion-based method for refining atomic models from cryo-EM density maps.
@@ 1 Introduction
Our ability to understand the molecular machinery of living organisms and design therapeutic compounds depends on our capability to observe the three-dimensional structures of protein and other biomolecular complexes. Experimental methods in structural biology such as X-ray crystallography, cryogenic electron microscopy (cryo-EM) and nuclear magnetic resonance (NMR) spectroscopy provide noisy and partial measurements from which the 3D structure of these biomolecules can be inferred. However, turning experimental observations into reliable 3D structural models is a challenging computational task. For many years, reconstruction algorithms were based on Maximum-A-Posteriori (MAP) estimation and often resorted to hand-crafted priors to compensate for the ill-posedness of the problem. State-of-the-art algorithms for cryo-EM reconstruction [56, 52] are instances of such “white-box” algorithms. These approaches can estimate the uncertainty of their answers and provide theoretical guarantees of correctness, but can only leverage explicitly defined regularizers and do not cope well with complex noise sources or missing data. Very recently, methods like Blush regularization in cryo-EM reconstruction [37] use a data-driven prior based on the noise2noise framework [42] to bypass the need for heuristic regularizers. However, the denoiser does not explicitly leverage the knowledge that proteins are made of atoms and cannot take advantage of a known amino acid sequence.
@@
Recently, supervised-learning approaches have emerged as an alternative to the MAP framework and some of them established a new empirical state of the art for certain tasks. The algorithm ModelAngelo [31], for example, can convert cryo-EM density maps into 3D atomic models with unprecedented accuracy. Typically, these supervised learning methods view the reconstruction problem as a regression task where a mapping between experimental measurements and atomic models needs to be learned. Some of these, like ModelAngelo, can even combine experimental data with sequence information by leveraging a pretrained protein language model [53]. However, these methods can only cope with a predefined type of input data. If additional information is available in a format that the model was not trained on (e.g., structural information about a fragment of the protein), or if the distribution of input data shifts at inference time (e.g., if the noise level changes due to modifications in the experimental protocol), a new model needs to be trained to properly cope with the new data.
@@
In the field of imaging, scenarios where an image or a 3D model must be inferred from corrupted and partial observations are known as “inverse problems”. To overcome the ill-posedness of these problems, regularizers were heuristically defined to inject hand-crafted priors and turn Maximum Likelihood Estimation (MLE) problems into MAP problems. In a similar fashion, machine learning-based methods were recently shown to outperform hand-crafted algorithms for a wide variety of tasks: denoising [74], inpainting [72], super-resolution [45], deblurring [49], monocular depth estimation [20], and camera calibration [36], among others. Recently, generative models were shown to be an effective way to inject data-driven priors into MAP problems, making inverse problems well-posed while replacing heuristic priors [6]. Among these generative methods, diffusion models gained popularity due to their powerful capabilities in the unconditional generation of images [19], videos [26], and 3D assets [51], and were recently leveraged to solve inverse problems in image space [59, 10].
The field of structural biology has also witnessed the application of diffusion models in protein structure modeling tasks [70, 1]. Building on recent progress in protein structure prediction [33] (sequence to structure) and fixed-backbone design [3] (structure to sequence), these diffusion models have opened the doors to de novo protein design [28, 70, 2, 73] (joint generation of structure and sequence). In particular, the recently released generative model Chroma [29] stands out in part thanks to its “programmable” framework, i.e. its ability to be conditioned on external hard or soft constraints. However, despite experimentally validated state-of-the-art results at unconditional generation and its programmable framework, Chroma has not yet been applied to inverse problems like atomic model building.
@@
Here, we introduce ADP-3D (Atomic Denoising Prior for 3D reconstruction), a framework to condition a diffusion model in protein space with any observations for which the measurement process can be physically modeled. Instead of using unadjusted Langevin dynamics for posterior sampling, our approach performs MAP estimation and leverages the data-driven prior learned by a diffusion model using the plug-n-play framework [67]. We demonstrate that our method handles a variety of external information: simulated cryo-EM density maps, amino acid sequence, partial 3D structure, and pairwise distances between amino acid residues, to refine a complete 3D atomic model of the protein. We show that our method outperforms posterior sampling baselines and, given a cryo-EM density map, can accurately refine incomplete atomic models provided by Modelangelo. In all our experiments, we use Chroma [29] as our pretrained generative model and highlight the importance of the diffusion-based prior. We therefore make the following contributions:
@@
  • We introduce a versatile framework, inspired from plug-n-play, to solve inverse problems in protein space with a pretrained diffusion model as a learned prior;
@@
  • We outperform existing posterior sampling methods at reconstructing full protein structures from partial structures;
@@
  • We show that a protein diffusion model can be guided to perform atomic model refinement in simulated cryo-EM density maps;
@@ 2 Related Work
Considerable progress has been made in leveraging diffusion models for protein structure generation. Lee et al. [41] demonstrated de novo protein generation using diffusion models over 2D distance matrices, requiring a post-processing step to produce 3D structures. Conditioned on secondary structure adjacency matrices, Anand and Achim [2] used a 3D point cloud representation to generate new structures. Trippe et al. [66] circumvented the need for conditioning information, demonstrated unconditional generation of proteins larger than 20 amino acid residues and introduced a novel conditional sampling algorithm to generate structures respecting a target motif. Wu et al. [71] built a generative model that directly operates on backbone internal coordinates (i.e., dihedral angles), thereby ensuring $\text{SE}(3)$-equivariance. Based on theoretical works extending diffusion models to Riemannian manifolds [16, 40], Yim et al. [73] introduced a principled formulation of diffusion models in $\text{SE}(3)$ and represented protein backbones as elements of $\text{SE}(3)^{N}$. In RFdiffusion, Watson et al. [70] experimentally designed the generated proteins and structurally validated them with cryo-EM. Recently, AlphaFold 3 [1] showed that a diffusion model operating on raw atom coordinates could be used as a tool to improve protein structure prediction. Diffusion models have been used to address specific tasks in protein space, like the generation of complementarity-determining region (CDR)-loops conditioned on non-CDR regions of antibody-antigen complexes [47], or sequence-to-structure prediction [32].
Out of the existing generative models for proteins, Chroma [29] has reported state-of-the-art “designability” metrics with a model that can be conditioned to generate proteins with desired properties (e.g., substructure motifs, symmetries). Despite its modular interface in which users can define their own conditioners, Chroma has not yet been used to guide the generative model with experimental measurements and solve real-world inverse problems. Here, we introduce a framework to efficiently condition a pretrained protein diffusion model and demonstrate the possibility of using cryo-EM maps as conditioning information.
@@
An inverse problem in image space can be defined by $\mathbf{y}=\Gamma(\mathbf{x})+\eta$ where $\mathbf{x}$ is an unknown image, $\mathbf{y}$ a measurement, $\Gamma$ a known operator and $\eta$ a noise vector of known distribution, potentially signal-dependent. The goal of posterior sampling is to sample $\mathbf{x}$ from the posterior $p(\mathbf{x}|\mathbf{y})$, the normalized product of the prior $p(\mathbf{x})$ and the likelihood $p(\mathbf{y}|\mathbf{x})$. Bora et al. [6] showed that generative models could be leveraged to implicitly represent a data-learned prior and solve compressed sensing problems in image space. Motivated by the success of diffusion models at unconditional generation [19], several works showed that score-based and denoising models could be used to solve linear inverse problems like super-resolution, deblurring, inpainting and colorization [43, 8, 55, 35, 46, 76], leading to results of unprecedented quality. Other methods leveraged the score learned by a diffusion model to solve inverse problems in medical imaging [60, 30, 9, 12, 11] and astronomy [61]. Finally, recent methods went beyond the scope of linear problems and used diffusion-based posterior sampling on nonlinear problems like JPEG restoration [59], phase retrieval and non-uniform deblurring [10]. Taking inspiration from these methods, we propose to leverage protein diffusion models to solve nonlinear inverse problems in protein space.
@@
Model building methods were first developed in X-Ray crystallography [13] and automated methods like Rosetta de-novo [68], PHENIX [44, 65] and MAINMAST [64] were later implemented for cryo-EM data. Although they constituted a milestone towards the automation of model building, obtained structures were often incomplete and needed refinement [58]. Supervised learning techniques were built for model building, relying on CNN and U-Net-based architectures [57, 75, 50]. EMBuild [25] was the first method to make use of sequence information and Modelangelo [31] established a new state of the art for automated de novo model building. Trained on 3,715 experimental paired datapoints, Modelangelo uses a GNN-based architecture and processes the sequence information with a pretrained language model [53]. Although fully-supervised methods outperform previous approaches, they still provide incomplete atomic models and cannot use a type of input data it was not trained with as additional information.
@@ 3 Background
Figure 1: Overview of ADP-3D. Our method turns partial and noisy measurements (the “conditioning information”) into a 3D structure by leveraging a pretrained diffusion model (here, Chroma [29]) and physics-based models of the measurement processes. Starting from a random structure $\mathbf{x}_{T}$, our method iterates between a denoising step and a data-matching step. The denoiser comes from the pretrained diffusion model. The data-matching step aims at maximizing the likelihood of the measurements.
3.1 Diffusion in Protein Space
We describe the atomic structure of a protein of $N$ amino acid residues by the 3D Cartesian coordinates $\mathbf{x}\in\mathbb{R}^{4N\times 3}$ of the four backbone heavy atoms (N, Cα, C, O) in each residue, the amino acid sequence $\mathbf{s}\in\{1,...,20\}^{N}$, and the side chain torsion angles for each amino acid $\mathbf{\chi}\in(-\pi,\pi]^{4N}$ (the conformation of the side chain can be factorized as up to four sequential rotations).
In Chroma [29], the joint distribution over all-atom structures is factorized as
@@ 3.1 Diffusion in Protein Space
Adding isotropic Gaussian noise to a backbone structure would rapidly destroy simple biophysical patterns that proteins always follow (e.g., the scaling law of the radius of gyration with the number of residues). Instead, Chroma uses a non-isotropic noising process as an inductive bias to alleviate the need for the model to learn these patterns from the data. The correlation of the noise is defined in such a way that a few structural properties are statistically preserved throughout the noising process. Specifically, the forward diffusion process is defined by the variance preserving stochastic process
@@
where $\beta_{t}$ is a time-dependent noising schedule and $\mathrm{d}\mathbf{w}$ is a standard Wiener process of dimension $\mathbb{R}^{4N\times 3}$. The matrix $\mathbf{R}\in\mathbb{R}^{4N\times 4N}$ is fixed and defined explicitly based on statistical considerations regarding the structure of proteins (see [29] and supplementary). Starting from $\mathbf{x}_{0}$ at $t=0$, a solution to this stochastic differential equation (SDE) at time $t$ is given by
@@
In practice, the SDE (4) is modified with a low-temperature sampling strategy in order to promote high-likelihoood states. The reverse-time low-temperature diffusion process is equilibrated with Langevin dynamics and solved numerically with the Euler-Maruyama method.
@@ 3.2 Half Quadratic Splitting and Plug-n-Play Framework
$\tilde{\mathbf{x}}_{k}$ $=\text{prox}_{g,\rho}(\mathbf{x}_{k})=\operatorname*{argmin}_{\tilde{\mathbf{x}}}~{}{g(\tilde{\mathbf{x}})+\frac{\rho}{2}\|\tilde{\mathbf{x}}-\mathbf{x}_{k}\|_{2}^{2}},$ (6)
$\mathbf{x}_{k+1}$ $=\text{prox}_{f,\rho}(\tilde{\mathbf{x}}_{k})=\operatorname*{argmin}_{\mathbf{x}}~{}{f(\mathbf{x})+\frac{\rho}{2}\|\mathbf{x}-\tilde{\mathbf{x}}_{k}\|_{2}^{2}},$
where prox are called “proximal operators” and $\rho>0$ is a user-defined proximal parameter.
If $f$ represents a negative log-likelihood over $\mathbf{x}$ and $g$ represents a negative log-prior, the above problem defines a Maximum-A-Posterior (MAP) problem. The key idea of the plug-and-play framework [67] is to notice that the first minimization problem in (6) is exactly a Gaussian denoising problem at noise level $\sigma=\sqrt{1/\rho}$ with the prior $\exp(-g(\mathbf{x}))$ in $\mathbf{x}$-space. This means that any Gaussian denoiser can be used to “plug in” a prior into a MAP problem.
@@ 4 Methods
In this section, we formulate our method, ADP-3D (Atomic Denoising Prior for 3D reconstruction), as a MAP estimation method in protein space and explain how the plug-n-play framework can be used to leverage the prior learned by a pretrained diffusion model. We then introduce our preconditioning strategy in the case of linear problems. Finally, we describe and model the measurement process in cryogenic electron microscopy. ADP-3D is described with pseudo-code in Algorithm 1.

### v2

@@ Abstract
The interaction of a protein with its environment can be understood and controlled via its 3D structure. Experimental methods for protein structure determination, such as X-ray crystallography or cryogenic electron microscopy, shed light on biological processes but introduce challenging inverse problems. Learning-based approaches have emerged as accurate and efficient methods to solve these inverse problems for 3D structure determination, but are specialized for a predefined type of measurement. Here, we introduce a versatile framework to turn biophysical measurements, such as cryo-EM density maps, into 3D atomic models. Our method combines a physics-based forward model of the measurement process with a pretrained generative model providing a task-agnostic, data-driven prior. Our method outperforms posterior sampling baselines on linear and non-linear inverse problems. In particular, it is the first diffusion-based method for refining atomic models from cryo-EM maps and building atomic models from sparse distance matrices.
@@ 1 Introduction
Experimental methods in structural biology such as X-ray crystallography, cryogenic electron microscopy (cryo-EM) and nuclear magnetic resonance (NMR) spectroscopy provide noisy and partial measurements from which the 3D structure of biomolecules can be inferred. This three-dimensional information is key for our understanding of the molecular machinery of living organisms, as well as for designing therapeutic compounds. However, turning experimental observations into reliable 3D structural models is a challenging computational task. For many years, reconstruction algorithms were based on Maximum-A-Posteriori (MAP) estimation and often resorted to hand-crafted priors to compensate for the ill-posedness of the problem. State-of-the-art algorithms for cryo-EM reconstruction [63, 59] are instances of such “white-box” algorithms. These approaches sometimes provide estimates for the uncertainty of their answers but can only leverage explicitly defined regularizers and do not cope well with complex noise sources or missing data.
@@
Recently, supervised-learning approaches have emerged as an alternative to the MAP framework and some of them established a new empirical state of the art for certain tasks, like model building [39]. Typically, these supervised learning methods view the reconstruction problem as a regression task where a mapping between experimental measurements and atomic models needs to be learned. Some of these, like ModelAngelo, can even combine experimental data with sequence information by leveraging a pretrained protein language model [60]. However, these methods must be trained on paired data (i.e., must be given input–output pairs) and can only cope with a predefined type of input. If additional information is available in a format that the model was not trained on (e.g., structural information about a fragment of the protein), or if the distribution of input data shifts at inference time (e.g., if the noise level changes due to modifications in the experimental protocol), a new model needs to be trained to properly cope with the new data.
@@
In the field of imaging, scenarios where an image or a 3D model must be inferred from corrupted and partial observations are known as “inverse problems”. To overcome the ill-posedness of these problems, regularizers were heuristically defined to inject hand-crafted priors and turn Maximum Likelihood Estimation (MLE) problems into MAP problems. In a similar fashion, machine learning-based methods were recently shown to outperform hand-crafted algorithms for a wide variety of tasks: denoising [86], inpainting [82], super-resolution [49], deblurring [53], monocular depth estimation [27], and camera calibration [42], among others. These methods, however, are equally limited by their need for paired data and their poor performance in the eventuality of a distribution shift.
In contrast, the MAP approach does not require paired data and can leverage the knowledge of the physics behind the problem through the definition of a likelihood function. As for the prior, generative models were shown to be effective tools to inject data-driven priors into MAP problems, making inverse problems well-posed while circumventing the need for heuristic priors [5]. Among these generative methods, diffusion models gained popularity due to their powerful capabilities in the unconditional generation of images [25], videos [34], and 3D assets [57], and were recently leveraged to solve inverse problems in image space [67, 12]. The field of structural biology has also witnessed the application of diffusion models in protein structure modeling tasks [79, 1]. The recently released generative model Chroma [37] stands out in part thanks to its “programmable” framework, i.e., its ability to be conditioned on external hard or soft constraints, but was never applied to structure determination problems like atomic model building.
@@
Here, we introduce ADP-3D (Atomic Denoising Prior for 3D reconstruction), a framework to condition a diffusion model in protein space with any observations for which the measurement process can be physically modeled. Instead of using unadjusted Langevin dynamics for posterior sampling, our approach performs MAP estimation and leverages the data-driven prior learned by a diffusion model using the plug-n-play framework [75], [88]. We demonstrate that our method handles a variety of external information: cryo-EM density maps, amino acid sequence, partial 3D structure, and pairwise distances between amino acid residues, to refine a complete 3D atomic model of the protein. We show that our method outperforms a posterior sampling baseline in average accuracy and, given a cryo-EM density map, can accurately refine incomplete atomic models provided by ModelAngelo. ADP-3D can leverage any protein diffusion model as a prior, which we demonstrate by showing results obtained with Chroma [37] and RFdiffusion [79]. We therefore make the following contributions:
@@
  • We introduce a versatile framework, inspired by plug-n-play, to solve inverse problems in protein space with a pretrained diffusion model as a learned prior;
@@
  • We outperform an existing posterior sampling method at reconstructing full protein structures from partial structures;
@@
  • We show that a protein diffusion model can be guided to perform atomic model refinement in simulated and experimental cryo-EM density maps;
@@ 2 Related Work
Considerable progress has been made in leveraging diffusion models for protein structure generation. While the first models could sample distance matrices [46], they were later improved to directly sample backbone structures represented by 3D point clouds [2, 74], backbone internal coordinates [81], 3D “frames” [84]. Recent methods are able to directly sample all-atom structures, including side chains [10]. In RFdiffusion, Watson et al. [79] experimentally designed the generated proteins and structurally validated them with cryo-EM. In Chroma, Ingraham et al. [37] introduced a “conditioning” framework to generate proteins with desired properties (e.g., substructure motifs, symmetries), but this framework was never used to enable protein structure determination from exprimental measurements. Recently, AlphaFold 3 [1] showed that a diffusion model operating on raw atom coordinates could be used as a tool to improve protein structure prediction. As generative models for proteins keep improving, leveraging them in the most impactful way becomes an increasingly important matter.
Here, we introduce a framework to efficiently condition a pretrained protein diffusion model and demonstrate the possibility of using cryo-EM maps as conditioning information. Most of our experiments are conducted using Chroma as a prior and we provide additional results using RFdiffusion in the supplements.
@@
An inverse problem in image space can be defined by $\mathbf{y}=\Gamma(\mathbf{x})+\eta$ where $\mathbf{x}$ is an unknown image, $\mathbf{y}$ a measurement, $\Gamma$ a known operator and $\eta$ a noise vector of known distribution, potentially signal-dependent. The goal of posterior sampling is to sample $\mathbf{x}$ from the posterior $p(\mathbf{x}|\mathbf{y})$, the normalized product of the prior $p(\mathbf{x})$ and the likelihood $p(\mathbf{y}|\mathbf{x})$. Bora et al. [5] showed that generative models could be leveraged to implicitly represent a data-learned prior and solve compressed sensing problems in image space. Motivated by the success of diffusion models at unconditional generation [25], several works showed that score-based and denoising models could be used to solve linear inverse problems like super-resolution, deblurring, inpainting and colorization [47, 9, 62, 41, 51, 88], leading to results of unprecedented quality. Other methods leveraged the score learned by a diffusion model to solve inverse problems in medical imaging [68, 38, 11, 14, 13] and astronomy [69]. Finally, recent methods went beyond the scope of linear problems and used diffusion-based posterior sampling on nonlinear problems like JPEG restoration [67], phase retrieval and non-uniform deblurring [12]. We refer to Daras et al. [19] for an in-depth survey of the methods leveraging diffusion models as priors in inverse problems. Taking inspiration from these methods, and in particular from DiffPIR [88], we propose to leverage protein diffusion models to solve nonlinear inverse problems in protein space.
@@
Model building methods were first developed in X-Ray crystallography [15] and automated methods like Rosetta de-novo [76], PHENIX [48, 73] and MAINMAST [72] were later implemented for cryo-EM data. Although they constituted a milestone towards the automation of model building, obtained structures were often incomplete and needed refinement [66]. Supervised learning techniques were applied to model building, relying on U-Net-based architectures [65, 87, 55], or combining a 3D transformer with a Hidden Markov Model [31]. EMBuild [33] was the first method to make use of sequence information and ModelAngelo [39] established a new state of the art for automated de novo model building. Trained on 3,715 experimental paired datapoints, ModelAngelo uses a GNN-based architecture and processes the sequence information with a pretrained language model [60]. Although fully-supervised methods outperform previous approaches, they still provide incomplete atomic models and cannot use a type of input data it was not trained with as additional information.
@@ 3 Background
Figure 1: Overview of ADP-3D. Our method turns partial and noisy measurements (the “conditioning information”) into a 3D structure by leveraging a pretrained diffusion model and physics-based models of the measurement processes. Starting from a random structure $\mathbf{x}_{T}$, our method iterates between a denoising step and a data-matching step. The denoiser comes from the pretrained diffusion model. The data-matching step aims at maximizing the likelihood of the measurements.
3.1 Diffusion in Protein Space with Chroma
In Chroma [37], the atomic structure of a protein of $N$ amino acid residues is represented by the 3D Cartesian coordinates $\mathbf{x}\in\mathbb{R}^{4N\times 3}$ of the four backbone heavy atoms (N, Cα, C, O) in each residue, the amino acid sequence $\mathbf{s}\in\{1,...,20\}^{N}$, and the side chain torsion angles for each amino acid $\mathbf{\chi}\in(-\pi,\pi]^{4N}$ (the conformation of the side chain can be factorized as up to four sequential rotations). The joint distribution over all-atom structures is factorized as
@@ 3.1 Diffusion in Protein Space with Chroma
Adding isotropic Gaussian noise to a backbone structure would rapidly destroy simple biophysical patterns that proteins always follow (e.g., the scaling law of the radius of gyration with the number of residues). Instead, Chroma uses a non-isotropic noising process as an inductive bias to alleviate the need for the model to learn these patterns from the data. The correlation of the noise is defined in such a way that a few structural properties are statistically preserved throughout the noising process. Specifically, the forward diffusion process is defined by the variance-preserving stochastic process
@@
where $\beta_{t}$ is a time-dependent noising schedule and $\mathrm{d}\mathbf{w}$ is a standard Wiener process of dimension $\mathbb{R}^{4N\times 3}$. The matrix $\mathbf{R}\in\mathbb{R}^{4N\times 4N}$ is fixed and defined explicitly based on statistical considerations regarding the structure of proteins (see Ingraham et al. [37] and supplements). Starting from $\mathbf{x}_{0}$ at $t=0$, a solution to this stochastic differential equation (SDE) at time $t$ is given by
@@ 3.2 Half Quadratic Splitting and Plug-n-Play Framework
[…]
where prox are called “proximal operators” and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}>0$ is a user-defined proximal parameter.
If $f$ represents a negative log-likelihood over $\mathbf{x}$ and $g$ represents a negative log-prior, the above problem defines a Maximum-A-Posterior (MAP) problem. The key idea of the plug-and-play framework [75] is to notice that the first minimization problem in (6) is exactly a Gaussian denoising problem at noise level $\sigma=\sqrt{1/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}}$ with the prior $\exp(-g(\mathbf{x}))$ in $\mathbf{x}$-space. This means that any Gaussian denoiser can be used to “plug in” a prior into a MAP problem.
@@ 4 Methods
In this section, we formulate our method, ADP-3D (Atomic Denoising Prior for 3D reconstruction), as a MAP estimation method in protein space and explain how the plug-n-play framework can be used to leverage the prior learned by a pretrained diffusion model. The method is described visually in Figure 1. We then introduce our preconditioning strategy in the case of linear problems. Finally, we describe and model the measurement process in cryogenic electron microscopy. ADP-3D is described with pseudo-code in Algorithm 1.